### v19

# All Zeros of the Riemann Zeta Function in the Critical Strip are Located on the Critical Line and are Simple

Abstract
In this paper we study the function β, for β. We derive a functional equation that relates and for all β, and we prove:
β’
That and the Riemann Zeta function have exactly the same zeros in the critical region β;
β’
The Riemann hypothesis, i.e., that all of the zeros of in are located on the critical line β; and that
β’
All the zeros of the Riemann Zeta function located on the critical line are simple.
Keywords: Riemann hypothesis, Fourier transforms, Schwarz
reflection principle, CauchyβRiemann equations
AMS Subject Classification: 11M06, 42A38
1 Introduction and Summary
The proof of the Riemann hypothesis is a problem that many mathematicians consider to be the most important problem of mathematics. Indeed, it is one of at most seven mathematics problems for which the Clay Institute has offered a million dollars for its solution. To this end, the .pdf publication [1] of Bombieri (see also [2]) presents an excellent summary β along with references, to papers and books, to connections with prime numbers, to Fermatβs last theorem, and to the work of authors who have shown that the first 1.5 billion zeros of the Zeta function listed with increasing imaginary parts are all simple β all of which are related to the mathematics of this subject. Similarly Wikipedia of the web [3] offers an excellent summary along with references about this subject. The magazine Nature recently published a related article about a discovery by Y. Zhang, of a conjecture (see [4]) on the spacing of prime numbers with increasing size. Physicists have also published results on the Riemann hypothesis: in [5], Meulens compares data about the Riemann hypothesis with solutions of two dimensional Navier Stokes equations, while others [6] have compared eigenvalues of selfβadjoint operators with zeros of the Riemann Zeta function. Several papers about solutions to the Riemann hypothesis have also appeared. To this end, the the papers of Violi [7], CoransonβBeadu, [8], GarciaβMorales, [9], and Chen [10] are similar to ours, in that their proof of the Riemann hypothesis are for functions that are different from the Zeta function, but which have the same zeros in as the Zeta function. We note the incredible amount of numerical work that has occurred, to find a zero of the zeta function in the critical strip but not on the critical line: see e.g. [11,12,13]. We also wish to mention the excellent work of [14], to establish that that all zeros of the Riemann zeta function are on the critical line, provided that all the zeros of the zeta function are simple. To this end, our present paper appears to be the only one which proves that all of the zeros of the Zeta function on the critical line are simple.
Castelvecchi, author of the article [4] makes the comment: βThe Riemann hypothesis will probably remain at the top of mathematicians wish lists for many years to come. Despite its importance, no attempts so far have made much progress.β
We wish of course to disagree with Castelvecchiβs comment at the end of the previous paragraph, since we believe that we have indeed proved the Riemann hypothesis in this selfβcontained paper, in which we accomplish the following:
1.
In Β§2, defining the function and showing in detail that has exactly the same zeros, in the critical strip, β, including multiplicity, as the Zeta function;
2.
Deriving a functional relation for in Β§2;
3.
Proving the positivity of for and the negativity of for in Β§2; where
(1.1) -
4.
Introducing the Schwarz reflection principle in Β§3, which the functions and of Equation (1.1) satisfy in β;
5.
Proving by contradiction, via use of results developed in Β§2 and in Β§4, that all of the zeros of G, (i.e., all of the zeros of the Zeta function) on the critical line are simple; and
6.
In Β§4, proving the Riemann hypothesis by contradiction, by use of results developed in Β§2, and in Β§3, of this paper, i.e., proving that (and the zeta function) have no zeros in the region β, where denotes the critical line, β.
Let β, and let β. In this paper we thus derive results about the function defined by the integral,
which is related to the wellβknown integral for the Riemann Zeta function, where
and where denotes the gamma function.
The operations of Schwarz reflection, the evaluation of and on important intervals of β, can be readily obtained from the Fourier transform representation of β, which is gotten from Equation (1.2) and which is defined for β, whereas an explicit Fourier transform of defined by Equation (1.3) for does not seem to be available.
2 Representations and related properties of
In this section, we introduce the Fourier transform representation of β, and we develop related properties of which we shall require in our proof in Β§4.
The function has many related properties, with the best known of these given by:
By setting and in (2.1), we get a Fourier integral representation of β, namely,
where β, β, and where is defined by
2.1 Properties of β, and
In this section we use the definition of given in Equation (1.2) and the identities of Equation (2.1) to derive a functional equation for β, and to derive additional properties of and β. We also show in detail, that and have exactly the same zeros in β, including multiplicity, that is positive for all β, and strictly decreasing as a function of β, for β, and we determine ranges of values of β, and their derivatives on the real line.
Let us next assign notations for the left and right half of the complex plane, the critical strip(s), and the critical line.
Definition 2.1
Let denote the left half of the complex plane, i.e., β, and let denote the right half of the complex plane. Let the critical strip be defined by β, and let the negative and positive critical strips be defined by β, and β. The critical line is defined by β, and the negative and positive critical lines are defined by and . Let and denote two distinct points of β, let denote the open line segment joining and β, and similarly for the half open, or closed, line segments joining and β.
2.2 Relevant Gamma function relations
We shall require the use of the following lemma:
Lemma 2.2
(i.) By replacing with in the duplication formula for the Gamma function, we get:
(ii.) Both and
are bounded by for all β, by
Equations (6.1.30) and (6.1.31) of [15]; and
(iii.) The function is an entire function;
is analytic in except for simple poles
at ().
Proof.Β
Item (i.) Is just Equation (6.1.18) of [15] with z replaced by
β;
Items (ii.) Follow from Equations (6.1.30) and (6.1.31)
of [15]; and
Item (iii.) Is just a restatement of a result found
in Chapter 16. of [15].
βΒ
2.3 Bounds on
The next lemma describes some asymptotic bounds on the function β, which are obtained by inspection of Equation (2.3).
Lemma 2.3
For any and for real, we have
Hence the integral is finite for any polynomial β.
βΒ
2.4 Analyticity definition of multiplicity
Definition 2.4
Let β, let denote an integer, and let be analytic in a neighborhood of β.
(a.) The function is said to have multiplicity at if there exists a finite number such that
(b.) If the multiplicity of at is β, and if β, then we shall more specifically say that is of exact multiplicity at β;
(c.) If is of exact multiplicity at β, then is said to be a zero (resp., a pole) of of multiplicity if (resp., if ). In particular, if β, (resp., if β,) then is said to be a simple zero (resp., a simple pole) of β;
(d.) We shall say that vanishes with total multiplicity on a line segment (or on β, or ), if there exist a total of distinct points (or on β, or ), with of multiplicity β, for β, such that
and we shall say that this total multiplicity on is exact if for all β.
2.5 Functional equations for and
An important identity of the Riemann Zeta function is the well known functional equation for , due to Riemann:
This functional equation for the Riemann Zeta function has many important uses, including, e.g., the analytic continuation of Zeta to all of β.
The function also possesses a functional equation which is given in Lemma 2.5 below, which plays a similar role as the functional equation for β. The functional equation for is gotten by substituting the rightβhandβside of the third equation of (2.1) into (2.5), and by use of Lemma 2.2:
Lemma 2.5
Let β, and let be defined as in (2.2). Then, a functional equation for the function , valid for all is:
This equation can also be written in a more compact form:
where
and where is nonβvanishing in β.
Proof.Β The identity (2.6) is valid for all since equation (2.5) is valid for all β, and since, except for simple poles, all of the functions multiplying the zeta function in the third equation of (2.1) are analytic in all of β. An alternate proof obtains by inspection of Equations (2.7) and (2.8). That is analytic and nonβvanishing in follows from Lemma 2.2 and by inspection of Equation (2.8).
βΒ
2.6 Zeros of and in
We prove here the and have the same zeros with the same multiplicity in β, and that these zeros are isolated.
Lemma 2.6
(i.) The functions and have exactly the same zeros in
β, including multiplicity; and
(ii.) All zeros of in are isolated.
Proof.Β (i.) By inspection if the third equation of (2.1) we get, if is a zero of of multiplicity β, then β, where the function is analytic and nonβvanishing in β, so that is also a zero of of multiplicity β. In addition, by taking the derivative of both sides of β, with nonβnegative integer β, we get,
Hence, if is a zero of of multiplicity β, then by applying induction to Equation (2.9), with respect to we conclude that the multiplicity of the zero of is also β;
(ii.) Suppose that there exists a cluster of zeros of in with a sub-sequence that has a limit point . If β, then would have to vanish, by Vitaliβs theorem. If is on the line β, then, since β, and since is analytic in β, it follows by use of the functional equation of , that where the point is now located on the line β, i.e., we are back to the previous case of the convergence of such a sub-sequence to a point on the interior of the right half plane, where is analytic and bounded, so that would again have to vanish identically in β.
βΒ
2.7 Definitions of and
Definition 2.7
Let and be defined as in Equation (2.2), and let us define as follows:
If for brevity, we write for and for and β, and for β, then Equation (1.2) yields the following definitions for and for β, where denotes a nonβnegative integer:
Let us define for any nonβnegative integer by
Then,
where these functions are readily shown to exist, by Lemma 2.3.
Lemma 2.8
Let the functions be defined as in Definition 2.7β. Then, for all β, and for all β, is analytic on the right half plane, and hence also in β. In particular given any β, is uniformly bounded in the region β.
Proof.Β This result follows directly by inspection of Equation (2.2) and Lemma 2.3. We omit the straightβforward proofs.
βΒ
2.8 Restricting the domain of
The following lemma restricts the domain of some of our identities:
Lemma 2.9
Let be defined by β, where the functions are defined in Definition 2.7. Then for all β, and moreover, is a strictly decreasing function of β for every fixed β.
Proof.Β We have
which shows is a strictly decreasing function of for all fixed β. By making the oneβtoβone transformation of in the above expressions for β, and then setting β, we get
Since is positive on β, and since is a strictly decreasing function of for all fixed β, we need only prove that for all β. To this end we have, by use of Taylor series expansions, that
Upon proceeding from the first to the second line of Equation (2.15) we used the following relations, which are valid for all β. The rightβhand side of Equation (2.15) then shows that for all β, i.e., that for all β.
βΒ
2.9 Inequalities for
The following lemma summarizes values of that have been established above.
Lemma 2.10
Let denote a nonβnegative integer. Then:
(i.) for all β;
(ii.) for all β; and
(iii.) for all β.
βΒ
Remark 2.11
1.
The functions β, β, β, and yield the correct sign of of these functions at β;
2.
Not so for the functions β, β, or ), which vanish identically on β; and
3.
If any of the functions of the previous item vanish on β, then it will be better to use β, β, β, or β, in place of them.
3 Schwarz reflection
We present the Schwarz reflection principle, which we define as follows:
Definition 3.1
Let be analytic in β, and real on β, for some β. Then can be continued analytically (i.e., reflected) across from to ) by means of the formula
Remark 3.2
The Schwarz reflection principle enables analytic continuation from to all of β. For example, if denotes a nonβnegative integer, so that the functions and are given for β, then by Lemma 2.10, and by Equation (2.8), is a nonβvanishing function of on β, while if β, then is a nonβvanishing function of β, changes sign as changes sign, while does not change sign as changes sign.
4 Proof of the Riemann Hypothesis
This section contains the proof of the following theorem:
Theorem 4.1
Let denote the critical region in the complex plane, and let β. Every zero of located on the critical line is a simple zero of β, and every zero of in is is located on the critical line β.
Proof.Β We shall prove this theorem by means of the proof by induction, and by contradiction, in two lemmas, which follow. We are able to carry out our proofs using and itβs derivatives instead of those of or β, see Remark 2.11.
The zeros of on are isolated, by Lemma 2.6. Hardy and Littlewood proved in [16] that the Zeta function, and hence β, has an infinite number of zeros on the critical line β. We begin this section by proving that all of the zeros of on are simple.
Lemma 4.2
Every zero of on the critical line is a simple zero of β.
Proof.Β Let us label all of the zeros of on in increasing order of imaginary parts as follows: β, where β.
We wish to prove that the case of a zero of of exact multiplicity greater than one is not allowed. Hence: Let be the proposition that the zeros of on are simple, for .
(I.) Proof that the case of a zero of of exact multiplicity greater than one is not allowed. To this end, let us assume the contrary, i.e., that each of and vanish with multiplicity β, where denotes a nonβnegative integer. Then vanishes on with total multiplicity β, or equivalently, vanishes on with multiplicity one, which is not allowed, by Lemma 2.10. This proves that the zero of of exact multiplicity greater than one is not allowed.
(I.2) Proof that if is true for all integers with β, for some β, then is true for β. If is true for β, then each of vanish with multiplicity one for all β.
Let us then assume for β, that both vanish with exact multiplicity β, where denotes a nonβnegative integer. Since vanishes at points on β, vanishes on with total multiplicity β. Equivalently, vanishes on with multiplicity one, but which is not allowed, by lemma 2.10.
This proves that the zero of of exact multiplicity greater than one is not allowed.
This completes the proof of Lemma 4.2.
βΒ
Lemma 4.3
If β, with β, and with β, then β.
Proof.Β If β, then for all β, by Lemma 2.10, so we shall assume, in the remainder of our proof which follows, that β.
We shall examine the vanishing of at points β, where for integer β, we have β, and where for each fixed β, with a positive integer, we shall assume that β. For simplicity of proof, we shall consider, for each fixed , the appearance of in the order β.
Let us state the induction hypothesis: Let denote a positive integer, and let be the proposition that for all β.
I.1: Proof that is true, for β.
Let us assume the opposite of the above induction hypothesis, i.e., that each of the two values, vanishes with multiplicity β, where denotes a nonβnegative integer. Then, we get, by reflection, that each of vanishes with multiplicity β, where denotes a nonβnegative integer. Hence if we assume in addition, that β, we get, by reflection, that vanishes on with total multiplicity β. Equivalently, vanishes with multiplicity one, but which is not allowed by Lemma 2.10.
On the other hand, if each of vanishes with multiplicity one, and if each of vanishes with multiplicity β, where denotes a nonβnegative integer, then vanishes on with total multiplicity β. Equivalently, then vanishes with multiplicity one, but which is also not allowed, by Lemma 2.10.
Hence the vanishing of with positive multiplicity is not allowed, regardless of whether or not vanishes. By the same argument, the vanishing of is not allowed for β, regardless of whether or not vanishes. Hence, the vanishing of is not allowed for all β.
Hence is true.
I.2: Proof that if is true, for β, where denotes a positive integer, then is true.
If is true, there are no points β, with for every fixed β, such that β. Hence we need to prove, that if is true, then is true, i.e., that β, for all β.
Let us assume the opposite, i.e., let us assume that both vanish with multiplicity β, where denotes a nonβnegative integer, and let us assume, in addition, that β. Then vanishes on with total multiplicity β, where denotes a nonβnegative integer. Equivalently, vanishes with multiplicity oneβ, but which is not allowed, by Lemma 2.10.
On the other hand, if vanishes, then it vanishes with multiplicity one, so that vanishes on with total multiplicity β, or equivalently, β, but which is not allowed, by Lemma 2.10. Hence we cannot allow the vanishing of with multiplicity β, where is the largest such value on β.
Next, the second largest, becomes the largest, and we can eliminate this case with the same proof that we used above to eliminate the case of β, and so on, until we have eliminated all of the vanishings of β, for β. Hence is true.
This completes the proof of Lemma 4.3.
βΒ
This completes the proof of Theorem 4.1.
βΒ
Acknowledgments:
The author is grateful to Drs. Paul Gauthier and Marc Stromberg for their corrections of the proofs related to Lemma 2.5.